OpenStudy (anonymous):

The x-coordinate of y–intercept of a line containing the points (6, 10) and (6, –10) is

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Can you find the equation of line?

OpenStudy (anonymous):

Can you find the equation of line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You have formula too. :P

OpenStudy (anonymous):

The equation of line passing through two points that are \((x_1,y_1)\) and \((x_2,y_2)\) is given by: \[\frac{y - y_1}{y_2 - y_1} = \frac{x-x_1}{x_2-x_1}\]

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

Find the equation of line dear. :)

OpenStudy (anonymous):

Great. :)

OpenStudy (anonymous):

ty:) @waterineyes

OpenStudy (anonymous):